A block of mass m kg is kept on a weighing machine in an elevator. If the elevator is retarding upward by a m/s², the reading of weighing machine is
(1) mg
(2) m(g-a)
(3) m(1-a/g)
(4) m(g+a)

Grade:11

5 Answers

Eshan
5 years ago
Dear student,

Elevator retarding upwards means that it is accelerating downwards with magnitude ‘a’. Hence block experience pseudo force ma upwards. Hence the reading of wieghing machine= m(g-a)
Mohd Suhail Ali
5 years ago
As lift goes upward net acc. is  g+a
But here a is retarding so -a should be taken so it will be g-a
Hence, N=m(g-a)  
In kgf the weighing machine reading= m(1-a/g) 
 
nishant narkar
5 years ago
The normal force tge acts on the blocl is m{g-a}
Therefore the weight shown will be m{g-a} please upvote
faz
5 years ago
here is ur solution for the question . kindly do upvte . this is just formula based and a easy problem
Srishti sharma
4 years ago
Since it is moving upward , therefore
N-mg=ma 
N= m(g+a)
And weighing machine measures Normal reaction not weight if a person . It measures weight in non inertial frame of reference where a = 0 .
So option no 4 is correct .
Thank you 😊